OpenStudy (anonymous):

solve 3/x + 5/x = -4

OpenStudy (zzr0ck3r):

multiply all terms by x 3+5=-4x then devide by -4 8/-4=x x=-2

Parth (parthkohli):

8/x = -4 -4x = 8 x = -2 SIMPLE RIGHT?

OpenStudy (anonymous):

yea

OpenStudy (zzr0ck3r):

only siple if you know how to do it

OpenStudy (zzr0ck3r):

simple*